Rectangle of a perimeter is 240cm. If the lenght of rectangle is decreased by 10% & breadth is increased by 20% the perimeter remain same(i.e. 240cm) find the length & breadth of rectangle
Answer by ankor@dixie-net.com(22740)

Rectangle of a perimeter is 240cm. If the length of rectangle is decreased by 10% & breadth is increased by 20% the perimeter remain same(i.e. 240cm) find the length & breadth of rectangle
:
"Rectangle of a perimeter is 240cm."
:
Let x = the original length; y = original width
:
2x + 2y = 240
Simplify divide equation by 2:
x + y = 120
or
y = (120-x); for substitution
:
"If the lenght of rectangle is decreased by 10% & breadth is increased by 20% the perimeter remain same(i.e. 240cm)"
New length = .9x
New width = 1.2y
:
2(.9x) + 2(1.2y) = 240
Simplify, divide by 2
.9x + 1.2y = 120
:
"find the length & breadth of rectangle"
:
Substitute(120-x) for y in the above equation, find x:
.9x + 1.2(120-x) = 120
.9x + 144 - 1.2x = 120
.9x - 1.2x = 120 - 144
-.3x = -24
x = -24/-.3
x = + 80 cm is the original length
:
Width: 120 - 80 = 40 cm is the original width
:
Find the new dimension:
.9 * 80 = 72 cm, new length
1.2 * 40 = 48 cm new width
:
Check the new perimeter
2(72) + 2(48) =
144 + 96 = 240